OpenStudy (anonymous):

Ryan left the science museum and drove south.three hours later gabriella driving 42km/h faster in and effort to catch up to him. After two hours gabriella finally caught up. find ryans average speed

OpenStudy (anonymous):

2*42 = 5*x So, x = 84/5 = 16.8km/h

OpenStudy (anonymous):

no the answer is 28 km/h i need the work let me know :)

OpenStudy (anonymous):

Oh....it is 42km/h FASTER. So, 2*(x+42) = 5*x => 2x + 84 = 5x => 3x = 84 => x = 84/3 => x = 28km/h